AM 166: Homework # 5 (due Thursday, March 22)

The book we refer to is Wackerly, Mendenhall & Scheaffer, Mathematical Statistics with Applications (6th Edition).

1. For some problems in this homework, you may want to use STATA to carry out the computation or draw the graph.

2. Solve problems 13.9, 13.13, 13.21, 13.30, 13.62 (a)(b).

3. Solve problems 15.2, 15.3, 15.18 (a), 15.19.

4. Solve problems 14.1, 14.8.